# Ruelle-Pollicott resonances of diffusive U(1)-invariant qubit circuits

### v1

@@ II.2 Truncation and Ruelle-Pollicott resonances
where $b,b^{\prime}\in\mathcal{P}^{(r)}$ (6), $m,m^{\prime}\in\{0,\dots,s-1\}$ and $k\in[0,2\pi)$. The RHS is expressed only in terms of local densities and the angled brackets denote the local Hilbert-Schmidt inner product $\left\langle a,b\right\rangle\coloneq\frac{1}{2^{N}}\mathrm{tr}\,a^{\dagger}b$. The written form is valid if $\delta r\leq s$. Importantly, the truncated propagator is block-diagonal with blocks indexed by the quasi-momentum $k$. More generally, the range of values of $j$ over which one has to sum in Eq. (10) depends on the spreading $\delta r$ of the circuit. For the derivation and further details see Appendix A.2.
The truncated propagator $\mathcal{U}^{(r)}_{k}$ is non-unitary and its eigenvalues lie inside the unit disc. One might think that its eigenvalues smoothly converge back to the unit circle in the unitary limit $r\to\infty$, but in many cases this is not true. If the local correlations decay exponentially one expects that the leading eigenvalue converges (or “freezes”) to some value inside the unit disc, which one interprets as the leading RP resonance $\lambda_{1}(k)$.
For the finite-sized truncated propagator, assuming it is diagonalizable, we can motivate that more explicitly. Let $\lambda_{i}^{(r)}$ be the eigenvalues of $\mathcal{U}_{k}^{(r)}$ (sorted by decreasing magnitude) and $\mathfrak{L}_{i}^{(r)}$/ $\mathfrak{R}_{i}^{(r)}$ its left/right eigenvectors. A generic infinite temperature correlation function of observables $A$, $B$ with good quasi-momentum $k$ can then be decomposed as
@@
$t\to\infty]{}\lambda_{1}^{t}\left\langle A^{\dagger},\mathfrak{R}_{1}^{(\infty)}\right\rangle\left\langle\mathfrak{L}_{1}^{(\infty)},B\right\rangle.$
@@
Therefore, if $A,B$ are not orthogonal to the eigenvectors associated with the leading RP resonance, their correlation function will decay as $\left\lvert\lambda_{1}\right\rvert^{t}$ at long times, which matches with the definition of an RP resonance in Eq. (2). Here on the LHS :-$\left\langle\bullet\right\rangle\coloneq\frac{1}{2^{N}}\mathrm{tr}\,(\bullet)⟨ ∙ ⟩ :$ (i.e., an infinite temperature average in a finite system) and, for clarity, $A,B$ are assumed to be traceless. If $A,B$ are not traceless (i.e., they overlap with identity, which is also the eigenvalue of $\mathcal{U}$), a similar argument can be made for the connected correlation function defined in Eq. (1). In all the following sections we shall discuss connected infinite temperature autocorrelation functions, we will usually shorten our language and simply say correlation functions. A similar procedure can also be done at finite temperature or chemical potential by introducing an appropriate inner product with a non-trivial measure.
@@
If a circuit has a conserved quantity, correlation functions of observables overlapping with a conserved quantity will converge to a non-zero value. In the RP resonance language, this would imply the presence of $\lambda_{i}=1$ in the decomposition in Eq. (11). Additionally, also all powers of the conserved quantity are conserved quantities, although typically non-local. While they cannot cause correlation functions of extensive (or local) observables to converge to the non-zero value in the thermodynamic limit, they can introduce plateaus in finite systems that decay as a power law in system size. We discuss this effect in Appendix D.
II.3 Example: Brickwall quantum circuits with no conserved quantities
@@
Let us demonstrate the above RP formalism with a simple example of a circuit with no conserved quantities. The main idea is to compare the exactly calculated correlation function in a finite system with the asymptotic decay obtained from the truncated propagator. In addition, we will show that the quasi-momentum dependence of $\lambda_{1}(k)$ can be non-monotonous, specifically, the fastest decay can happen at nonzero $k$.
@@
We take a brickwall circuit with nearest-neighbor 2-qubit gates, all being the same. The gate is picked randomly according to the unitary Haar measure. In Fig. 2b we show the leading RP resonance $\left\lvert\lambda_{1}(k)\right\rvert$ for two realizations (cases) of a random local gate. For details on numerical methods see Appendix B.1. The circuit’s Floquet propagator can be written as
@@
where $V_{i,j}$ are local gates, i.e., 2-site unitary operators acting on sites $i$ and $j$. All $V$ are taken to be the same. In both cases we see that the $k$ dependence of $\lambda_{1}(k)$ is highly non-trivial and depends strongly on the choice of $V$. We show $k$ dependence only for $k\in[0,\pi]$, since $\left\lvert\lambda_{1}(k)\right\rvert=\left\lvert\lambda_{1}(-k)\right\rvert$. This is a consequence of the fact that $\left(\mathcal{U}^{(r)}_{k}\right)^{*}=\mathcal{U}^{(r)}_{-k}$, which follows from the form in Eq. (10). One must only show that the $\left\langle\mathcal{S}^{sj+m}(b),\mathcal{U}\left(\mathcal{S}^{m^{\prime}}(b^{\prime})\right)\right\rangle$ is a real number, which quickly follows from the hermiticity of the Pauli basis, $b=b^{\dagger},b^{\prime}=b^{\prime\dagger}$.
Fig. 2c shows the decay of correlation functions of extensive observables with $k=0$ and $\pi$. For details about numerical calculation of correlation functions see Appendix B.2. Their long-time decay rate matches well with the leading RP resonance determined from the truncated propagator. If the observables do not have good quasi-momentum $k$, for instance strictly local observables, their correlation function expands over the whole spectrum of eigenvalues with different momenta $k$. The long time decay is then governed by the biggest $\lambda_{1}(k)$. It is often assumed that this happens at $k=0$, i.e., for translationally invariant observables, but as we see in Fig. 2b for case ii, this is not always the case. There the gap seems to be the smallest around $k\approx 2$.
In the following sections we shall focus on the main subject of our work, namely, study how a conserved quantity changes the RP spectrum and what can we extract from it. As mentioned in Sec. I.1, RP resonances can also be studied in the presence of conservation laws by introducing Lindbladian dissipation Mori (2024), though we find the truncated propagator more convenient; we briefly discuss the Lindblad approach in Appendix C.
Refer to caption(a)                                                             Refer to caption(b)
Figure 3: Circuits with conserved magnetization. (a) Circuit diagram with magnetization preserving 3-qubit gate $V$ (the shown circuit has $s=3$ and $\delta r=6$), whose structure with non-zero blocks is shown on the left. (b) The leading RP eigenvalue of the truncated propagator with support $r$ for 3 choices of $V$ (chosen according to the Haar measure). The same cases will be used throughout the paper.
III The leading Ruelle-Pollicott resonance
@@
Therefore, our system, in addition to a translational symmetry, also has a U(1) symmetry. A natural way to achieve that would be to take a brickwall circuit with a 2-qubit gate that conserves magnetization. However, if the gate is the same everywhere such circuits are all special – they are integrable Žnidaric et al. (2024); Žnidarič (2025a), including for non-brickwall geometries Paletta et al. (2025). Second possibility that also preserves translational symmetry would be to take a brickwall configuration with two different magnetization-conserving gates, one in each brickwall layer, though it turns out that such circuits have slow (possibly non-diffusive) thermalization Jonay et al. (2024).
Because we want good diffusive thermalization, we therefore consider one of the next simplest options, which are circuits with 3-site gates. The gates are arranged in a brickwall-like geometry with $s=3$ site translational invariance, see Fig. 3a. The Floquet propagator is given by the following expression
@@
with periodic boundary conditions (indices are taken modulo $N$). 3-site gates $V$ are taken the same everywhere, so that we have translational symmetry. In order for the circuit to conserve magnetization, each local gate must conserve it as well. One can thus quickly deduce that the local gate $V$ must have 4 independent blocks, which are shown on the diagram in Fig. 3a. We choose the matrix elements in each block independently according to the unitary Haar measure in the corresponding block, which results in a unitary $V$.
@@
The considered circuit spreads observables by $\delta r=6$ in one period (see Fig. 3) and, importantly, we have $\delta r>s$. Therefore, Eq. (10) for the matrix elements of the truncated propagator is not applicable. While one can use the more general formula derived in Appendix A.2, this approach gives access only up to $r=6$, since the evaluation of matrix elements is numerically performed on observables of support $r+2\delta r$. A substantial improvement can be made by noticing that successive layers are translations of the previous layers, that is $\mathcal{S}(U_{i})=U_{i+1}$. This space-time symmetry Duh and Žnidarič (2024) allows us to redefine the first layer (and a translation to the left) to be the “new” Floquet propagator for a three times shorter period, leading to less spreading $\delta r=2$ and allowing us to use Eq. (10). For details see Appendix A.3.
@@
The leading eigenvalue of the truncated propagator (and thus the leading RP resonance) for $3$ different realizations (cases with different gate $V$) is shown in Fig. 3b. We will use the same cases throughout this paper. Its $k$ dependence has two regimes, $k$ close to 00 (small $k$) and $k$ far from zero (around $k=\pi$). At $k=0$, the leading eigenvalue is exactly $1$ and its eigenvector is the local density corresponding to magnetization (i.e., the conserved quantity). Eigenvalues at small quasi-momenta have smooth dependence on $k$ and converge fast with $r$. We will show that they are related to transport and contain information about the spin diffusion constant in Sec. III.1. Eigenvalues far from $k=0$ converge slower and show more varied behavior, similarly to what we have seen in cases without symmetries (cf. Fig. 2b), including non-smooth dependence due to collisions with smaller eigenvalues. In Sec. III.2 we will numerically verify that a finite gap there indeed signals exponential decay of corresponding observables that are unrelated to transport.
Refer to captionRefer to captionRefer to caption(a)(a.i)(a.ii)
Refer to captionRefer to captionRefer to caption(b)(b.i)(b.ii)
Figure 4: Transport of magnetization in circuits case 1 (a) and case 2 (b). In both cases we show the leading RP resonance ($r=11$) and the fit to ${\rm e}^{-Dk^{2}}$ (insets i), and the domain-wall evolution (TEBD, $N=204$ with $\mu=10^{-3}$ and $\chi=256$) at times with equal spacing $\Delta t=22$ up to $t=200$ (main plots) with the associated transferred magnetization and the diffusion constant fit (insets ii).
III.1 Small quasi-momenta and magnetization transport
@@
In generic, i.e., chaotic, quantum systems with 1 conserved quantity, one typically expects diffusive transport Bertini et al. (2021). Diffusive transport can be probed by considering the correlation function of the local density of the conserved quantity, also called the dynamic structure factor, for which one obtains
@@
where $D$ is the model-dependent (spin) diffusion constant. We have adjusted our notation to $\sigma^{z}(x,t)$ meaning the $z$-Pauli matrix acting on space coordinate $x$ and propagated to time $t$. For clarity, we can also consider space to be continuous with $\Delta x=1$ corresponding to three sites of our model (i.e., we average our system over one translational period). Eq. (15) follows from the system obeying the diffusion equation on hydrodynamic scales. The diffusion equation can be derived by assuming linear relationship between expectation values of the spin current and the gradient of the density of the conserved quantity Forster (1990).
@@
The correlation function of magnetization in momentum space therefore decays exponentially, with the decay time diverging as $1/(Dk^{2})$ in the long wavelength limit $k\to 0$. This is a standard indicator of diffusion Forster (1990). Because there are evidently exponentially decaying correlation functions with their rate having quadratic $k$ dependence this implies that there must exist a leading RP resonance
@@
This can alternatively be interpreted as an RP continuum governing the decay of the local magnetization correlation function. We discuss other potential RP continuums greater detail in Sec. IV.2.
Since the RP resonance predicted in Eq. (16) converges to $1$ as $k\to 0$, it must be the leading RP resonance for small $k$. It is, therefore, precisely the RP resonance observed for small $k$ in Fig. 3b. This is confirmed numerically by fitting a function ${\rm e}^{-Dk^{z}}$ to $\lambda_{1}$ for small $k$, where $z$ is usually referred to as the transport dynamical exponent. This is shown in Fig. 4, insets (i). The obtained values of $z$ match the predicted diffusive $z=2$ up to accessible precision. One can also check the corresponding eigenvectors, which, according to the calculation in Eq. (16), should be the local density of magnetization. We have confirmed that this is the case at big enough $r$. At smaller $r$ the corresponding eigenvector includes other components. It would be interesting to extend this to a different type of transport, perhaps a non-diffusive one ($z\neq 2$) with a one-parameter scaling relation $x^{z}\sim t$. We expect the quasi-momentum dependence of $\lambda_{1}(k)$ for small $k$ to also hold information about $z$ in that case, although the exact form of $\lambda_{1}(k)$ depends not just on $z$, but also on the scaling function of the local correlation function of the conserved quantity.
The fit also extracts the spin diffusion constant $D$ of the model. In order to verify that this is indeed the correct value, we compare it to the diffusion constant obtained by a completely different method. Namely, from the evolution of an initial weakly-polarized domain wall state Ljubotina et al. (2017). We prepare the system in the state with the following density matrix
@@
for small $\mu$, and evolve it with $U$. Diffusion equation dictates the transferred magnetization $\Delta\left\langle\sigma^{z}\right\rangle_{\rho}$ should be (at long times in our system)
$\Delta\left\langle\sigma^{z}\right\rangle_{\rho}(t)\coloneq\sum_{j=N/2+1}^{N}\left\langle\sigma_{j}^{z}(t)\right\rangle_{\rho}=2\mu\sqrt{\frac{D}{\pi}t},$ (19)
where $\left\langle\bullet\right\rangle_{\rho}\coloneq\mathrm{tr}\,(\rho\,\bullet)$ denotes the expectation value in state $\rho$. The dynamics can be simulated using time-evolving block decimation (TEBD) Vidal (2003, 2004), for details see Appendix B.3. Fitting a curve scaling as $t^{1/z}$ to $\Delta\left\langle\sigma^{z}\right\rangle$ one can numerically confirm diffusion and extract the spin diffusion constant $D$.

### v2

@@ 2.2 Truncation and Ruelle-Pollicott resonances
where $b,b^{\prime}\in\mathcal{P}^{(r)}$ (6), $m,m^{\prime}\in\{0,\dots,s-1\}$ and $k\in(-\pi,\pi]$. The RHS is expressed only in terms of local densities and the angled brackets denote the local Hilbert-Schmidt inner product, $\left\langle a,b\right\rangle\coloneq\frac{1}{2^{N}}\mathrm{tr}\,a^{\dagger}b$. Importantly, the truncated propagator is block-diagonal with blocks indexed by the quasi-momentum $k$. The written form is valid if $\delta r\leq s$. More generally, the range of values of $j$ over which one has to sum in Eq. (10) depends on the spreading $\delta r$ of the circuit. For the derivation and further details see Appendix A.2.
The truncated propagator $\mathcal{U}^{(r)}_{k}$ is non-unitary and its eigenvalues lie inside the unit circle. One might think that its eigenvalues smoothly converge back to the unit circle in the unitary limit $r\to\infty$, but in many cases this is not true. If the local correlations decay exponentially one expects that the leading eigenvalue converges (or “freezes”) to some value inside the unit circle, which one interprets as the leading RP resonance $\lambda_{1}(k)$.
For the finite-$r$ truncated propagator of an infinite circuit $N\to\infty$, assuming it is diagonalizable, we can illustrate that more explicitly. Let $\lambda_{i}^{(r)}$ be the eigenvalues of $\mathcal{U}_{k}^{(r)}$ (sorted by decreasing magnitude) and $\mathfrak{L}_{i}^{(r)}$/ $\mathfrak{R}_{i}^{(r)}$ its left/right eigenvectors. A generic infinite-temperature correlation function of observables $A$, $B$ with good quasi-momentum $k$ can then be decomposed as
@@
$\xrightarrow[t\to\infty]{}\lambda_{1}^{t}\left\langle A^{\dagger},\mathfrak{R}_{1}^{(\infty)}\right\rangle\left\langle\mathfrak{L}_{1}^{(\infty)},B\right\rangle.$
@@
Therefore, provided that $A,B$ have a finite overlap with the eigenvectors associated with the isolated leading eigenvalue, their correlation function will decay as $\left\lvert\lambda_{1}\right\rvert^{t}$ at long times, which matches with the definition of an RP resonance in Eq. (2). Here on the LHS $\left\langle\bullet\right\rangle\coloneq\frac{1}{2^{N}}\mathrm{tr}\,(\bullet)$ (i.e., an infinite-temperature average in a finite system) and, for clarity, $A,B$ are assumed to be traceless. If $A,B$ are not traceless (i.e., they overlap with the identity, which is also the eigenvalue of $\mathcal{U}$), a similar argument can be made for the connected correlation function defined in Eq. (1). In all the following sections we shall discuss connected infinite-temperature autocorrelation functions, we will usually shorten our language and simply say correlation functions. A similar procedure can also be done at finite temperature or chemical potential by introducing an appropriate inner product with a non-trivial measure.
@@
If a circuit has a conserved quantity, correlation functions of observables overlapping with the conserved quantity will converge to a non-zero value. In the RP resonance language, this would imply the presence of $\lambda_{i}=1$ in the decomposition in Eq. (11). Additionally, also all powers of the conserved quantity are conserved quantities, although typically non-local. While they cannot cause correlation functions of extensive (or local) observables to converge to a non-zero value in the thermodynamic limit, they can introduce plateaus in finite systems that decay as a power law in system size. We discuss this effect in Appendix D.
2.3 Example: Brickwall quantum circuits with no conserved quantities
Refer to caption (b)                  Refer to caption(c)
Figure 2: Circuits with no conserved quantities. (a) Diagram of operator spreading in a brickwall circuit with the same gate $V$ acting between all nearest neighbors (the shown circuit has $s=2$, $\delta r=2$). (b) The leading eigenvalue of the propagator truncated to the space of extensive observables with support $r$ for two realizations of a circuit with a different gate $V$ chosen randomly according to the unitary Haar measure. (c) Infinite temperature autocorrelation functions of extensive observables AA (4) in the circuit with gate i from (b). The local density $a$ is either $\sigma^{z}$, or chosen randomly according to the Gaussian unitary ensemble (GUE) [27] with support $r=2$. Dashed lines show the RP prediction $\propto\left\lvert\lambda_{1}\right\rvert^{t}$, while full curves are an exact calculation in a circuit with $N=32$ qubits; the $\sigma^{z},k=0$ case is multiplied by $15$ for better presentation.
@@ 2.3 Example: Brickwall quantum circuits with no conserved quantities
Let us demonstrate the above RP formalism with a simple example of a circuit with no conserved quantities. The main idea is to compare the exactly calculated correlation function in a large finite system with the asymptotic decay obtained from the truncated propagator. In addition, we will show that the quasi-momentum dependence of $\lambda_{1}(k)$ can be non-monotonous, specifically, the slowest decay can happen at nonzero $k$.
@@
We take a brickwall circuit with nearest-neighbor 2-qubit gates, all being the same. The gate that is repeated in the circuit is picked randomly according to the unitary Haar measure. In Fig. 2b, we show the leading RP resonance $\left\lvert\lambda_{1}(k)\right\rvert$ for two realizations. For details on numerical methods see Appendix B.1. The circuit’s Floquet propagator can be written as
@@
where $V_{i,j}$ are local gates, i.e., 2-site unitary operators acting on sites $i$ and $j$. All $V$ are taken to be the same. In both realizations of a circuit we see that the $k$ dependence of $\lambda_{1}(k)$ is highly non-trivial and depends strongly on the choice of $V$. A simillarly complicated behavior was also observed for the kicked Ising model in Ref. [53]. We show the $k$ dependence only for $k\in[0,\pi]$, since $\left\lvert\lambda_{1}(k)\right\rvert=\left\lvert\lambda_{1}(-k)\right\rvert$. This is a consequence of the fact that $\left(\mathcal{U}^{(r)}_{k}\right)^{*}=\mathcal{U}^{(r)}_{-k}$, which follows from the form in Eq. (10). One must only show that $\left\langle\mathcal{S}^{sj+m}(b),\mathcal{U}\left(\mathcal{S}^{m^{\prime}}(b^{\prime})\right)\right\rangle$ is a real number, which quickly follows from the hermiticity of the Pauli basis, $b=b^{\dagger},b^{\prime}=b^{\prime\dagger}$.
Fig. 2c shows the decay of correlation functions of extensive observables with $k=0$ and $\pi$. For details about numerical calculation of correlation functions see Appendix B.2. Their long-time decay rate matches well with the leading RP resonance determined from the truncated propagator. If the observables do not have good quasi-momentum $k$, for instance strictly local observables, their correlation function expands over the whole spectrum of eigenvalues with different momenta $k$. The long time decay is then governed by the biggest $\lambda_{1}(k)$. It is often assumed that this happens at $k=0$, i.e., for translationally invariant observables, but as we see in Fig. 2b for the circuit with gate ii, this is not always the case. There the gap seems to be the smallest around $k\approx 0.6\pi$.
In the following sections we shall focus on the main subject of our work, namely, study how a conserved quantity changes the RP spectrum and what can we extract from it. As mentioned in Sec. 1.1, RP resonances can also be studied by introducing Lindbladian dissipation [29], though we find the truncated propagator more convenient; we briefly discuss the Lindblad approach in Appendix C.
3 The leading Ruelle-Pollicott resonance
@@
Therefore, our system, in addition to a translational symmetry, also has a U(1) symmetry. A natural way to achieve that would be to take a brickwall circuit with a 2-qubit gate that conserves magnetization. However, if the gate is the same everywhere, such circuits are all special – they are integrable [52, 50], including for non-brickwall geometries [32]. The second possibility that also preserves translational symmetry would be to take a brickwall configuration with two different magnetization-conserving gates, one in each brickwall layer, though it turns out that such circuits have slow (possibly non-diffusive) thermalization [15].
(a)                                     Refer to caption(b)
Figure 3: Circuits with conserved magnetization. (a) Circuit diagram with magnetization-conserving 3-qubit gate $V$ (the shown circuit has $s=3$ and $\delta r=6$), whose block structure is shown on the left. (b) The leading RP eigenvalue of the truncated propagator with support $r$ for 3 choices of $V$ (chosen according to the Haar measure). The same gates will be used throughout the paper.
Because we want good diffusive thermalization, we therefore consider one of the next simplest options, which are qubit circuits with 3-site gates. The gates are arranged in a 3-layer brickwall-like geometry with $s=3$ site translational invariance, see Fig. 3a. The Floquet propagator is given by the following expression
@@
with periodic boundary conditions (indices are taken modulo $N$). 3-site gates $V$ are taken the same everywhere, so that we have translational symmetry. In order for the circuit to conserve magnetization, the local gate $V$ must conserve it as well. One can thus quickly deduce that the local gate $V$ must have 4 independent blocks, which are shown on the diagram in Fig. 3a. To get a typical representative of such circuits, we choose the matrix elements in each block independently according to the unitary Haar measure in the corresponding block, which results in a unitary $V$.
@@
The considered circuit spreads observables by $\delta r=6$ in one period (see Fig. 3) and, importantly, we have $\delta r>s$. Therefore, Eq. (10) for the matrix elements of the truncated propagator is not applicable. While one can use the more general formula derived in Appendix A.2, this approach gives access only up to $r=6$, since the evaluation of matrix elements is performed numerically by the evolution of observables on $r+2\delta r$ sites. A substantial improvement can be made by noticing that successive layers are translations of the previous layer, that is $\mathcal{S}(U_{i})=U_{i+1}$. This space-time symmetry [6] allows us to redefine the first layer (and a translation to the left) to be the “new” Floquet propagator for a three times shorter period, leading to less spreading $\delta r=2$, allowing us to use Eq. (10) and thereby enabling us to reach about $r=13$. For details see Appendix A.3.
@@
The leading eigenvalue of the truncated propagator (and thus the leading RP resonance) for $3$ different realizations with different gate $V$ is shown in Fig. 3b. We will use the same gates throughout this paper. Its $k$ dependence has two regimes, $k$ close to 0 (small $k$) and $k$ far from zero (around $k=\pi$). At $k=0$, the leading eigenvalue is exactly $1$ and its eigenvector is the local density corresponding to magnetization (i.e., the conserved quantity $a=\sigma^{z}$). Eigenvalues at small quasi-momenta have smooth dependence on $k$ and converge fast with $r$. We will show that they are related to transport and contain information about the spin diffusion constant in Sec. 3.1. Eigenvalues far from $k=0$ converge slower and show more varied behavior, similarly to what we have seen in circuits without symmetries (cf. Fig. 2b), including non-smooth dependence due to collisions with smaller eigenvalues. In Sec. 3.2 we will numerically verify that a finite gap there indeed signals exponential decay of corresponding observables that are unrelated to transport.
3.1 Small quasi-momenta and magnetization transport
@@
In generic, i.e., chaotic, quantum systems with 1 conserved quantity, one typically expects diffusive transport [1]. Diffusive transport can be probed by looking at the correlation function of the local density of the conserved quantity, also called the dynamic structure factor. At sufficiently large $t$, it is expected to obey the diffusion equation [7], which gives
@@
where $D$ is the model-dependent (spin) diffusion constant. We have adjusted our notation to $\sigma^{z}(x,t)$ meaning the $z$-Pauli matrix acting on space coordinate $x$ and propagated to time $t$. In Eq. (15), we consider space to be continuous, with unit distance $\Delta x=1$ corresponding to three sites (i.e., the translational period) of the model.
@@
The correlation function of magnetization in momentum space therefore decays exponentially, with the decay time diverging as $1/(Dk^{2})$ in the long wavelength limit $k\to 0$. This is a standard indicator of diffusion [7]. Because there are evidently exponentially decaying correlation functions with their rate having quadratic $k$ dependence, this implies that there must exist an RP resonance
@@
This can alternatively be interpreted as an RP continuum governing the decay of the local magnetization correlation function. We discuss other potential RP continuums in greater detail in Sec. 4.2.
Refer to captionRefer to captionRefer to caption(a)(a.i)(a.ii)
Refer to captionRefer to captionRefer to caption(b)(b.i)(b.ii)
Figure 4: Transport of magnetization in the circuit with gate 1 (a) and gate 2 (b). In both instances we show the leading RP resonance ($r=11$) and the fit to ${\rm e}^{-Dk^{z}}$ (insets i), and the domain-wall evolution (TEBD, $N=354$ with $\mu=10^{-3}$ and $\chi=256$) at times with equal spacing $\Delta t=55$ up to $t=500$ (main plots) with the associated transferred magnetization and the diffusion constant fit (insets ii).
Since the RP resonance predicted in Eq. (17) converges to $1$ as $k\to 0$, it must be the leading RP resonance for small $k$. It is, therefore, precisely the RP resonance observed for small $k$ in Fig. 3b. This is confirmed numerically by fitting a function ${\rm e}^{-Dk^{z}}$ to $\lambda_{1}$ for small $k$, where $z$ is usually referred to as the transport dynamical exponent. This is shown in Fig. 4, insets (i). The obtained values of $z$ match the predicted diffusive $z=2$ up to accessible precision. One can also check the corresponding eigenvectors, which, according to the calculation in Eq. (16), should overlap with the local density of magnetization. We have confirmed that this is indeed the case. It would be interesting to extend this to a type of non-diffusive transport ($z\neq 2$) with a one-parameter scaling relation $x^{z}\sim t$. We expect the quasi-momentum dependence of $\lambda_{1}(k)$ for small $k$ to also hold information about $z$ in that case, although the exact form of $\lambda_{1}(k)$ depends not just on $z$, but also on the scaling function of the local correlation function of the conserved quantity.
The fit also extracts the spin diffusion constant $D$ of the model. In order to verify that this is indeed the correct value, we compare it to the diffusion constant obtained from a completely different method, namely, from the evolution of a weakly-polarized domain-wall state [21]. We prepare the system in the state with the following density matrix
@@
for small $\mu$, and evolve it with $U$. The diffusion equation dictates that the transferred magnetization $\Delta\left\langle\sigma^{z}\right\rangle_{\rho}$ should be (after long enough time)
$\Delta\left\langle\sigma^{z}\right\rangle_{\rho}(t)\coloneq\sum_{j=N/2+1}^{N}\left\langle\sigma_{j}^{z}(t)-\sigma_{j}^{z}(0)\right\rangle_{\rho}=2\mu\sqrt{\frac{D}{\pi}t},$ (19)
where $\left\langle\bullet\right\rangle_{\rho}\coloneq\mathrm{tr}\,(\rho\,\bullet)$ denotes the expectation value in the state $\rho$. The dynamics can be simulated using time-evolving block decimation (TEBD) [44, 45], for details see Appendix B.3. Fitting a curve scaling as $t^{1/z}$ to $\Delta\left\langle\sigma^{z}\right\rangle$ one can numerically confirm diffusion and extract the spin diffusion constant $D$.
Both methods of determining $D$ are shown in Fig. 4. In the transferred magnetization from the domain wall (bottom left (ii) insets), we again see $z$ matching the expected $z=2$. The diffusion constants determined from different approaches also match up to accessible precision of about $1\%$. We have thus shown the leading RP resonance for small $k$ truly does contain information about transport. A natural next question, which we leave for future work, is how competitive is the extraction of $D$ through RP resonances compared to the current state-of-the-art.
3.2 Quasi-momenta far from zero and correlation function decay